A sample of CO2 gas at 25 ºC has a volume of 15 L and exerts a pressuer of 15.6 psi. How many grams of gas are in the sample?




1
Expert's answer
2021-04-07T11:33:00-0400
"PV=nRT=(m\/M)RT"

"15.6 psi=1.08*10^5 K"

"m=MPV\/(RT)=44*1.08*10^5*0.015\/(8.314*298)=28.77g"